# Beyond Prompt Engineering: Robust Behavior Control in LLMs via Steering Target Atoms

### v1

@@ Abstract
Precise control over language model generation is vital for ensuring both safety and reliability. Although prompt engineering and steering are commonly used to intervene in model behaviors, the vast number of parameters in models often results in highly intertwined internal representations. This interdependency can limit control precision and sometimes lead to unintended side effects. Recent research has explored the use of sparse autoencoders (SAE) to disentangle knowledge in high-dimensional spaces for steering. However, these applications have been limited to toy tasks owing to the nontrivial issue of locating “atomic knowledge components”. In this paper, we propose Steering Target Atoms (STA), a novel method that isolates and manipulates disentangled knowledge components to enhance safety. Comprehensive experiments demonstrate the effectiveness of our approach. Further analysis reveals that steering exhibits superior robustness and flexibility, particularly in adversarial scenarios. We also apply the steering strategy to the large reasoning model, confirming its effectiveness in precise reasoning control.
@@
Mengru Wang1,211footnotemark: 1,  Ziwen Xu1thanks:   Equal Contribution.,  Shengyu Mao1, Shumin Deng3,  Zhaopeng Tu2,  Huajun Chen1,  Ningyu Zhang1thanks:   Corresponding Author. 1Zhejiang University,  2Tencent AI Lab, 3National University of Singapore, NUS-NCS Joint Lab, Singapore
@@ 1 Introduction
Steering vs. Prompting. We further conduct a comprehensive analysis to compare steering and prompting (§5). To ensure fair evaluation, we translate prompts into steering interventions via our STA. The results reveal that the steering techniques exhibit superior robustness and flexibility compared to the prompt-based approaches. From the perspective of previous observation Todd et al. (2024), both prompting and steering manipulate model behaviors by influencing internal computations. However, steering enables finer-grained control by directly modifying neuron activations in LLMs during forward propagation, whereas prompting relies on the model’s ability to infer behavior from input text. This may make steering more precise and robust, particularly when input signals degrade across layers Merullo et al. (2024); Dong et al. (2021), while prompting remains more intuitive and accessible.
@@ 4.3 Controlling Analysis
5 Controlling LLMs: Steering or Prompting?
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, We gratefully acknowledge the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ Appendix A SAE Training
Due to computational resource constraints, we do not train sparse autoencoders (SAEs) from scratch. Instead, we utilize the publicly available Gemma Scope Lieberum et al. (2024b) for our experiments.
@@ B.1.1 Prompting Strategy
We adopt the manual prompt from in Xie et al. (2023) as $\text{Prompt}_{\text{hand}}$ for safe defense:
@@
We apply the following template to automatically generate the prompt, denoted as $\text{Prompt}_{\text{auto}}$, for the specific task:

### v2

@@ Abstract
Precise control over language model generation is vital for ensuring both safety and reliability. Although prompt engineering and steering are commonly used to intervene in model behaviors, the vast number of parameters in models often results in highly intertwined internal representations. This interdependency can limit control precision and sometimes lead to unintended side effects. Recent research has explored the use of sparse autoencoders (SAE) to disentangle knowledge in high-dimensional spaces for steering. However, these applications have been limited to toy tasks owing to the nontrivial issue of locating “atomic knowledge components”. In this paper, we propose Steering Target Atoms (STA), a novel method that isolates and manipulates disentangled knowledge components to enhance safety. Comprehensive experiments demonstrate the effectiveness of our approach. Further analysis reveals that steering exhibits superior robustness and flexibility, particularly in adversarial scenarios. We also apply the steering strategy to the large reasoning model, confirming its effectiveness in precise reasoning control [^1] Code is available at https://github.com/zjunlp/steer-target-atoms..
@@
Mengru Wang1,211footnotemark: 1,  Ziwen Xu1thanks:   Equal Contribution.,  Shengyu Mao1, Shumin Deng3,  Zhaopeng Tu2,  Huajun Chen1,  Ningyu Zhang1thanks:   Corresponding Author. 1Zhejiang University,  2Tencent, 3National University of Singapore, NUS-NCS Joint Lab, Singapore
@@ 1 Introduction
Steering Vectors vs. Prompt Engineering. We further conduct a comprehensive analysis to compare steering and prompting (§5). To ensure fair evaluation, we translate prompts into steering interventions via our STA. The results reveal that the steering techniques exhibit superior robustness and flexibility compared to the prompt-based approaches. From the perspective of previous observation Todd et al. (2024), both prompt engineering and steering vectors manipulate model behaviors by influencing internal computations. However, steering vectors enable finer-grained control by directly modifying neuron activations in LLMs during forward propagation, whereas prompting relies on the model’s ability to infer behavior from input text. This may make steering more precise and robust, particularly when input signals degrade across layers Merullo et al. (2024); Dong et al. (2021), while prompting remains more intuitive and accessible.
@@ 4.3 Controlling Analysis
5 Controlling LLMs: Steering Vectors or Prompt Engineering?
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00). We gratefully acknowledge the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ Appendix A SAE Training
Due to computational resource constraints, we do not train sparse autoencoders (SAEs) from scratch. Instead, we utilize the publicly available Gemma Scope Lieberum et al. (2024b) and LlamaScope (He et al., 2024) for our experiments.
@@ B.1.1 Prompting Strategy
We adopt the manual prompt from in Xie et al. (2023) as $\text{Prompt}_{{hand}}$ for safe defense:
@@
We apply the following template to automatically generate the prompt, denoted as $\text{Prompt}_{{auto}}$, for the specific task: